OpenStudy (anonymous):

the floor of a rectangular shed is 7 ft longer than it is wide and has an area of 120 square ft. what are its dimensions?

OpenStudy (anonymous):

Let s = width l= length l=(s+7) ls=120 (s+7)s=120 solve for s

OpenStudy (anonymous):

and then sub the value of s into l=(s+7) to find l

OpenStudy (anonymous):

Thats the part I got stuck on. I got s^2+7s=120. I can't figure out how to get s from there

OpenStudy (anonymous):

s^2+7s-120=0 (s+15)(s-8)=0 s=-15, s=8

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

no problem